OpenStudy (anonymous):

z^2+9=10z I rewrote it as z^2-10z+9=0 but I don't understand what to do next can anyone please help me?????? thanks for nothing folks

OpenStudy (ajprincess):

What do u get when u multiply 16 by 81 @leslieJampolsky?

OpenStudy (anonymous):

1296

OpenStudy (ajprincess):

Can u tell me two factors of 1296 such that their sum is 72?

OpenStudy (anonymous):

36X36=1296 and 36+36=72

OpenStudy (ajprincess):

yup u r right:)

OpenStudy (anonymous):

I got to : 16x+36x+36x+81 but I don't know what to do next

OpenStudy (ajprincess):

Try to factorise 16x^2+36x and 36x+81. What do u get when u factorise them @leslieJampolsky?

OpenStudy (anonymous):

4x(4x+9) +9(4x+9)

OpenStudy (anonymous):

4x(4x+9)1(36+81) but I think that is wrong

OpenStudy (anonymous):

to uzumakhi: so to simplify even more would it read: x(4x+9)(4x+9)????

OpenStudy (anonymous):

no its (4x+9)^2

OpenStudy (ajprincess):

16x^2+36x=4x(4x+9) 36x+81=9(4x+9) 16x^2+36x+36x+81=4x(4x+9)+9(4x+9) =(4x+9)(4x+9) =(4x+9)^2 Getting this @leslieJampolsky?

OpenStudy (anonymous):

Notice how \(4x\cdot4x = 16x^2\) and \(9\cdot9 = 81\) and \(2(9+4x) = 72x\)

OpenStudy (anonymous):

thank you all so much

OpenStudy (ajprincess):

yw:)